Nissley Wedding Fantasy Corporation makes very elaborate wedding cakes to order. The owner of the company has provided the following data concerning the activity rates in its activity-based costing system:

Activity Cost Pools Activity Rate
Size-related $ 1.14 per guest
Complexity-related $ 28.48 per tier
Order-related $ 74.98 per order

The measure of activity for the size-related activity cost pool is the number of planned guests at the wedding reception. The greater the number of guests, the larger the cake. The measure of complexity is the number of tiers in the cake. The activity measure for the order-related cost pool is the number of orders. (Each wedding involves one order.) The activity rates include the costs of raw ingredients such as flour, sugar, eggs, and shortening. The activity rates do not include the costs of purchased decorations such as miniature statues and wedding bells, which are accounted for separately.

Data concerning two recent orders appear below:

Tijerina Wedding Twersky Wedding
Number of reception guests 44 132
Number of tiers on the cake 3 5
Cost of purchased decorations for cake $ 29.60 $ 54.90

Assuming that all of the costs listed above are avoidable costs in the event that an order is turned down, what amount would the company have to charge for the Tijerina wedding cake to just break even?
Answer #1

Answer

Amount the company would have to charge for the Tijerina wedding cake to just break even = $ 240.18

Calculations:

Cost
reception guests $                              50.16
tiers on the cake $                              85.44
Order cost $                              74.98
decorations for cake $                              29.60
Total $                            240.18

Note:

Cost
reception guests =1.14 * 44
tiers on the cake =28.48 * 3
Order cost =1 * 74.98
decorations for cake 29.6

in case of any doubt, please comment.